OpenStudy (princessstyles-horan):

2) Hunter is pulling on his little red wagon to the west with a force of 125N. Luke doesn’t want him to take the wagon so he is pulling in the opposite direction with a force of 75N. If the wagon has a mass of 15kg, what is the wagon’s acceleration? * Your answer

OpenStudy (shamim):

U hv to usr the equation F=ma

OpenStudy (shamim):

Acceleration a=?

OpenStudy (shamim):

Both F nd m r given

OpenStudy (shamim):

U hv 2 forces

OpenStudy (shamim):

Forces r opposite in direction

OpenStudy (shamim):

So resultant force F=125-75=?

OpenStudy (shamim):

This resultant force acting along west

OpenStudy (shamim):

So acceleration will b along west

OpenStudy (michele_laino):

hint: on the block are acting two forces, which have the same direction, and opposite orientation:  so the net force has the subsequent magnitude: \(R=F-P=125-75=...newtons\) and the requested acceleration, has the subsequent magnitude \(a\): \[a = \frac{R}{m} = \frac{R}{{15}} = ...m/{\sec ^2}\]